### v1

# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

Abstract
The widespread application of Deep Learning across diverse domains hinges critically on the quality and composition of training datasets. However, the common lack of disclosure regarding their usage raises significant privacy and copyright concerns. Dataset auditing techniques, which aim to determine if a specific dataset was used to train a given suspicious model, provide promising solutions to addressing these transparency gaps. While prior work has developed various auditing methods, their resilience against dedicated adversarial attacks remains largely unexplored. To bridge the gap, this paper initiates a comprehensive study evaluating dataset auditing from an adversarial perspective. We start with introducing a novel taxonomy, classifying existing methods based on their reliance on internal features (IF) (inherent to the data) versus external features (EF) (artificially introduced for auditing). Subsequently, we formulate two primary attack types: evasion attacks, designed to conceal the use of a dataset, and forgery attacks, intending to falsely implicate an unused dataset. Building on the understanding of existing methods and attack objectives, we further propose systematic attack strategies: decoupling, removal, and detection for evasion; adversarial example-based methods for forgery. These formulations and strategies lead to our new benchmark, DATABench, comprising 17 evasion attacks, 5 forgery attacks, and 9 representative auditing methods. Extensive evaluations using DATABench reveal that none of the evaluated auditing methods are sufficiently robust or distinctive under adversarial settings. These findings underscore the urgent need for developing a more secure and reliable dataset auditing method capable of withstanding sophisticated adversarial manipulation. Code is available at https://github.com/shaoshuo-ss/DATABench.
I Introduction
The rapid advancement of Deep Learning (DL) has propelled it into widespread applications across various domains and real-world systems, such as computer vision [32], artificial intelligence (AI) chatbots [28], and self-driving vehicles [10]. Currently, DL models primarily make predictions based on their knowledge learned from a dataset. As such, its success highly relies on and is inseparable from high-quality datasets [67, 45, 50]. As a critical resource and fundamental component of DL, datasets largely determine the performance of the models trained on them. However, developers of DL models and applications usually do not disclose the training datasets they use. This lack of transparency gives rise to potential privacy and copyright concerns. For instance, training data may include unauthorized private information. Commercially deployed models may have been trained using open-source copyrighted datasets that prohibit commercial use. Additionally, many existing data protection regulations, such as GDPR, have granted data owners the right to be informed of their data usage. Consequently, auditing the datasets used in DL models has become a critical issue for safeguarding data privacy and copyright [41, 18].
Figure 1: The general pipeline of dataset auditing. In the preparation stage, the data owner collects and publishes a dataset. In the exploitation stage, an adversary may copy or steal the dataset and train a model for unauthorized purposes. Finally, in the auditing stage, the data owner audits the suspicious model and verifies whether it is trained on the protected dataset.
Dataset auditing [^1] Existing literature may use different terminologies, such as dataset ownership verification [49], dataset copyright auditing [18], or data-use auditing [41]. In this paper, we standardize the term as dataset auditing in DL. in DL [49, 18, 42] refers to determining whether a specific dataset is used to train a suspicious model. Typically, as shown in Figure 1, there are three stages involved in dataset auditing [53]. In the first stage, the data owner collects some data from various resources. Subsequently, he processes the dataset ($e.g.$, data annotation) to make it suitable for usage or release. Then, in the second stage, an adversary may get access to the dataset through copying or stealing and utilize it to train a DL model without authorization. Third, if a third-party DL model (dubbed a ‘suspicious model’) is suspected of using a protected dataset unauthorizedly, dataset auditing techniques will be adopted to verify it. However, there exist several challenges to achieving a reliable and accurate dataset auditing. (1) It is hard to uncover the traces of a dataset only from the parameters, intermediate results, or even black-box access to a trained DL model. These revealed clues may not be directly associated with the data. (2) The adversary can use various different model architectures and training techniques to produce the model. The diversity and the training-agnostic situation make it difficult to achieve robust and distinctive dataset auditing. (3) The adversary may intentionally process the data or model to compromise the auditing procedure, challenging the robustness of data auditing solutions. As such, achieving dataset auditing is a non-trivial issue in practice [18, 71, 13].
Although some existing works have proposed promising solutions to dataset auditing [79, 29, 60], there is still a significant gap between how the robustness and distinctiveness of existing works are evaluated (often in controlled environments) and how they perform in real-world practice. Existing dataset auditing research primarily considers naive operations and techniques in DL under constrained settings, such as training with models of different architectures, fine-tuning, and pruning, as potential adversarial vectors [49, 18, 31]. These works lack the consideration of a ‘clever’ adversary who may design specified and adaptive attacks against dataset auditing. It is challenging to conduct an exhaustive evaluation due to the complexity of training a model in DL and the diversity of attack surfaces. Consequently, the following critical question remains unanswered: are existing dataset auditing methods reliable under adversarial settings?
In this paper, we initiate the first comprehensive study on evaluating dataset auditing in DL from an adversarial perspective. Our contributions are as follows.
  • Contribution 1: A Novel Taxonomy of Existing Methods.
We first present the formal definition of dataset auditing. We then revisit existing auditing methods and introduce a novel taxonomy of them. Beyond simply classifying existing methods according to the capabilities ($e.g.$, white/black-box) and operations ($e.g.$, intrusive/non-intrusive) of the data owner, we split the features of a protected dataset into internal features (IF) and external features (EF), and correspondingly classify auditing methods into IF-based and EF-based methods. IF refers to the feature that inherently exists in the data, and the model learns from IF to gain utility. IF-based methods analyze the characteristics of IFs ($e.g.$, overfitting) to achieve auditing. In contrast, EF is the artificial feature introduced by dataset auditing methods and can lead the model to have defender/owner-specified behaviors ($e.g.$, misclassification). EF-based methods achieve auditing by attempting to trigger these behaviors. This IF/EF distinction provides deeper insight into the underlying mechanisms of auditing methods and critically facilitates the systematic design of attacks, which previous taxonomies did not readily support.
  • Contribution 2: Unified Formulations of Potential Attacks against Dataset Auditing.
We formally formulate two fundamental adversarial objectives against dataset auditing, including evasion attacks and forgery attacks. Evasion attacks aim to conceal the usage of a protected dataset. This type of attack maliciously leads the auditing method to incorrectly determine that the victim dataset was not used for training. Conversely, forgery attacks seek to falsely attribute dataset usage. They deceive the auditing method into concluding that an adversary-designated unused dataset was part of the training dataset of an independent victim model. Furthermore, we systematically identify three key phases where adversaries can intervene: the pre-processing, training, and post-training phases. This comprehensive formulation provides a structured understanding of the attack surface in dataset auditing.
  • Contribution 3: Systematic Attack Strategies.
We propose systematic strategies for crafting effective attacks against dataset auditing, building upon our IF/EF taxonomy and attack formulations. For evasion, we introduce three core strategies according to targeted feature type: (1) Decoupling: Weakening the model’s reliance on or the detectability of IFs from the training dataset. (2) Removal: Eliminating the artificially introduced EFs or their corresponding designated behaviors. (3) Detection: Identifying inputs likely to trigger EF-based auditing mechanisms and circumventing the response. For forgery, we propose a unified strategy leveraging adversarial example techniques to mimic the specific signals (IF-based confidence responses or EF-based designated behaviors) expected by the target auditing method. These attack strategies provide concrete guidance for designing attacks and comprehensively evaluating dataset auditing.
  • Contribution 4: Dataset Auditing Attack Benchmark and Extensive Evaluation on Existing Methods.
We develop the Dataset AudiTing Attack Benchmark, DATABench, based on the proposed attack formulations and strategies. Specifically, we implement 17 evasion attacks and 5 forgery attacks, and evaluate 9 representative dataset auditing methods on the proposed benchmark. The empirical results reveal that none of the evaluated auditing methods exhibit sufficient robustness or distinctiveness under adversarial settings. We surprisingly find that even simple attacks can significantly compromise the effectiveness of existing methods. The results highlight an urgent need for the development of a more reliable dataset auditing method. Based on our findings, we also point out several potential future directions that can provide guidance to develop more reliable dataset auditing methods.
DATABench also provides an extensible open-source toolbox to facilitate future research and development in this crucial area. This toolbox is designed to streamline the process for researchers to easily develop new dataset auditing techniques, implement additional attacks, and systematically evaluate their proposed methods against the comprehensive and standardized attack suite provided within our unified benchmark.
II Background
II-A Deep Learning
Deep Learning (DL) has become the most famous and widely-used technique to develop an artificial intelligence (AI) application during the last decade [32, 67, 33], especially with the growth of large foundation models [28]. DL primarily leverages deep neural networks (DNNs) as the core model architecture to achieve a wide range of functions.
During the development of a DL model $\mathcal{M}$, the model developer needs to first collect a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{|\mathcal{D}|}$. The dataset $\mathcal{D}$ is related to the developer-specified DL task. The dataset consists of various input samples $\bm{x}_{i}$ and their corresponding label $\bm{y}_{i}$. Subsequently, the developer designs a loss function $\mathcal{L}(\cdot)$ and optimizes the model $\mathcal{M}$ according to the following optimization problem:
^$\widehat{\mathcal{M}}=\arg\min_{\mathcal{M}}\sum_{i=1}^{k}\mathcal{L}(\mathcal{M}(\bm{x}_{i}),\bm{y}_{i}).$ (1)
Currently, most of the DL techniques and models are data-centric and their developments highly rely on a high-quality dataset [67]. This reliance, however, raises significant concerns regarding data privacy and copyright. For instance, many DL models are trained on large-scale datasets that may contain sensitive or proprietary information, leading to potential violations of user privacy and intellectual property rights [6, 18]. Consequently, data protection and building high-quality datasets are equally important in practice.
II-B Data Protection
Data protection has long been a critical aspect of information management [62]. In the last several decades, researchers have developed many different techniques to achieve various requirements of data protection, as follows.
Encryption. Encryption-based techniques aim to transform the sensitive data into unreadable ciphertext, and any unauthorized party without the secret key cannot access the data [5, 25]. Advancements in encryption-based techniques like homomorphic encryption (HE) [64, 14] now allow computations on encrypted data, making it promising to be adopted in DL systems. However, encryption-based methods still suffer from the limitations of excessive latency and resource demands.
Differential Privacy. Differential privacy (DP) ensures that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output [22]. This is typically achieved by injecting controlled noise into data, queries, or model training processes [22, 1]. While DP provides strong, mathematically provable privacy guarantees, its application in DL introduces significant challenges. For instance, the added noise often degrades model utility. The trade-off between privacy and utility makes DP a powerful but costly solution for privacy-preserving DL [1, 85].
Traditional Data Watermarking. Traditional data watermarking techniques help track ownership and usage of data by embedding hidden markers ($i.e.$, watermark) into the data [43, 87]. The data owner can extract the watermark from the data and prove their ownership. Multiple solutions to embedding a watermark into data have been proposed, such as embedding watermarks into the frequency domain or utilizing DL models [8, 4]. Recently, some works have also explored using data watermarking to mark AI-generated content and models [86, 44, 69]. However, traditional data watermarking techniques may not be feasible for dataset auditing since the watermark in the data can hardly be inherited by the DL models and their owners will not disclose their training details (especially their training data).
Unlearnable Data. Unlearnable data techniques aim to add perturbations to the data so that the DL model cannot learn from it during training [39, 81]. The model trained on the unlearnable dataset will have a low utility. However, in the scenario of protecting a public dataset where we intend to prevent unauthorized usage instead of making it completely useless, these techniques may not be feasible.
In general, previous data protection techniques either have a significant impact on the data and model utility ($e.g.$, HE, DP, and unlearnable data) or are not feasible ($e.g.$, traditional data watermarking) in DL scenarios [49, 31]. Consequently, dataset auditing, as a technique that has negligible influence on the data utility and can protect the privacy and copyright of data owners, has become an emerging yet important and promising solution to data protection [47, 18].
III Dataset Auditing in Deep Learning
In this section, we present the formal definition, our novel taxonomy, and the objectives of dataset auditing methods. In this paper, we focus on dataset auditing for image classification models, since most existing methods work on them, and they usually cannot be directly adapted to other modalities and tasks. Further discussion about dataset auditing for generative models can be found in Section VI-B.
III-A Towards a Formal Definition of Dataset Auditing
Before defining dataset auditing, we hereby first introduce the general threat model of dataset auditing in DL [49, 18]. Specifically, there are two parties involved in the threat model, the data owner and the adversary [^2] Some existing literature [56, 76, 18] also introduces a trustworthy third-party auditor (or verifier) in their threat model to conduct the audit. However, since the auditor does not affect the technical design of dataset auditing methods, we exclude it to avoid potential confusion.. The data owner wants to collect, store, and publish a DL dataset. However, there exists the risk that an adversary may unlawfully copy or steal the dataset and train his own model for unauthorized usage. Such a misbehavior may compromise the privacy or intellectual property rights of the data owner. As such, the data owner seeks an effective dataset auditing method to confirm whether the suspicious model of the adversary is trained on his dataset. Now, we formally define dataset auditing as follows.
Definition 1 (Dataset Auditing).
Given a dataset $\mathcal{D}$, a suspicious model $\mathcal{M}(\cdot)$, and some auxiliary information $\mathcal{I}$, a dataset auditing method $\mathcal{A}(\cdot)$ outputs $1$ if the model is (partly) trained on the dataset $\mathcal{D}$, and 0 otherwise:
$\mathcal{A}(\mathcal{D},\mathcal{M},\mathcal{I})=\left\{\begin{aligned} 1,\ &\mathcal{M}\ \text{is trained on}\ \mathcal{D}\\ 0,\ &\mathcal{M}\ \text{is not trained on}\ \mathcal{D}\\ \end{aligned}\right..$ (2)
The auxiliary information $\mathcal{I}$ in Definition 1 varies depending on the specific dataset auditing method. Common forms of $\mathcal{I}$ include auxiliary datasets used for auditing [77], specific trigger patterns [49], and watermark information [41].
Notably, if the number of samples in $\mathcal{D}$ is $1$ ($i.e.$, $|\mathcal{D}|=1$), the definition of dataset auditing is similar to membership inference [6, 65, 74]. As such, existing membership inference techniques can be easily adapted to achieve dataset auditing, which will be further discussed in Section III-B.
Main Pipeline of Dataset Auditing. In general, there are three stages in a typical pipeline of dataset auditing.
  1. 1.
    Preparation Stage. The data owner collects the data and organizes it into a raw dataset $\mathcal{D}$. Then, before publishing the dataset, the data owner can also apply a processing function $\mathcal{P}(\cdot)$: $\widehat{\mathcal{D}}=\mathcal{P}(\mathcal{D},\mathcal{I}),$ where $\mathcal{I}$ is the auxiliary information needed in the processing function. $\mathcal{P}(\cdot)$ may be used to add specific perturbations or watermarks to the dataset, or just keep $\widehat{\mathcal{D}}=\mathcal{D}$. Then, the data owner publishes the dataset $\widehat{\mathcal{D}}$.
  2. 2.
    Exploitation Stage. After the data owner publishes the dataset $\widehat{\mathcal{D}}$, an adversary may steal or copy $\widehat{\mathcal{D}}$ to train his own model $\mathcal{M}$ for unauthorized purposes. He can adopt any training techniques and conduct certain operations to intentionally compromise the dataset auditing mechanism.
  3. 3.
    Auditing Stage. If the data owner finds a model [^3] Following prior works [18, 49, 41, 53], we do not focus on how to identify these ‘suspicious models’, which may be discovered through non-technical ways. For instance, if a model’s task closely aligns with that of a dataset, the model may be regarded as a ‘suspicious model’. that is suspected of unauthorizedly using the dataset $\widehat{\mathcal{D}}$, he adopts a dataset auditing method $\mathcal{A}$ to judge whether the misbehavior exists. If $\mathcal{A}(\widehat{\mathcal{D}},\mathcal{M},\mathcal{I})$ outputs $1$, the adversary can be accused of unauthorizedly using the dataset and infringing the copyright or privacy.
Assumptions of Data Owner. The data owner intends to publish a dataset and audit any unauthorized usage of it. The capabilities of the data owner are as follows.
  • In the preparation stage, the data owner has full control of this stage and the dataset. He can determine what preprocessing steps $\mathcal{P}(\cdot)$ to apply to the dataset and which portions of the data to release.
  • In the exploitation stage, the data owner has no knowledge of the operations of the adversary. He also cannot interfere with the training of the adversary’s model.
  • In the auditing stage, the data owner only has access to the adversary’s trained model.
According to the operation and capability of the data owner, existing dataset auditing methods can be further classified according to the following simple taxonomies [18].
Definition 2 (Intrusive and Non-intrusive Dataset Auditing).
In the preparation stage, if the data owner does not need to alter the raw dataset, $i.e.$, $\mathcal{D}=\widehat{\mathcal{D}}$, the dataset auditing method is a non-intrusive method. Otherwise, it is an intrusive method.
Definition 3 (White-box and Black-box Dataset Auditing).
In the auditing stage, if $\mathcal{A}$ only needs API access to the suspicious model ($i.e.$, only get the outputs of given inputs), it is defined as a black-box method. Otherwise, if additional information about the suspicious model is needed ($e.g.$, the parameters or the intermediate results during inference), the dataset auditing method $\mathcal{A}$ is a white-box method.
In this paper, we pay more attention to black-box dataset auditing since it is more practical in real-world scenarios.
III-B Dataset Auditing Taxonomy
Existing works [18] primarily introduce simple taxonomies such as Definitions 2 or 3. However, these taxonomies offer limited aid in understanding the underlying insights of dataset auditing methods and designing effective attack strategies to evaluate their reliability. As such, we propose a new taxonomy based on the origin of the features exploited during the auditing process. Specifically, we split the features of the data into internal and external features and classify existing dataset auditing methods into the following two main categories.
Internal Feature (IF)-based Dataset Auditing. This class of methods relies on the internal feature of the dataset. Internal features refer to features that naturally exist within the data itself. Learning from these features builds up the utility of the model on the primitive task. The core insight of IF-based dataset auditing is to exploit the characteristics of the internal features ($e.g.$, overfitting) during or after training.
Dataset auditing methods based on membership inference attacks (MIA) [77, 35, 55] fall into this category. MIA aims to determine whether a given data instance was included in a model’s training dataset. Given a trained model and a specific data point, MIA outputs a confidence score that reflects the likelihood that the instance was part of the training set. Although originally proposed as a privacy attack, MIA can be readily adapted to the task of dataset auditing by aggregating the confidence scores across all the samples in a dataset ($e.g.$, calculating the average confidence score).
Additionally, some existing works explore auxiliary mechanisms to enhance the effectiveness of MIA in the context of dataset auditing. Some studies propose Dataset Inference (DI) [60, 23], which treats the dataset as a whole and utilizes the prediction margin to conduct dataset auditing. [59] also explores aggregating results of different MIAs. Data-use Auditing (DUA) [41] adds perturbations to craft two datasets that are significantly different from each other in the feature space. DUA achieves dataset auditing by testing whether the confidence of being a member on one dataset is significantly larger than the other. These mechanisms make the IF-based dataset auditing more feasible in practice.
External Feature (EF)-based Dataset Auditing. In contrast, EF-based dataset auditing methods introduce external or artificial features into the dataset. The model trained on the protected dataset with external features will have identifiable behavior that is hardly learned from other datasets. This class of methods achieves dataset auditing by verifying whether the suspicious model can trigger the designated behavior.
Existing methods have explored different ‘identifiable behaviors’ for auditing. Radioactive Data [71, 3] induces distributional shifts in feature space. Misclassification on specific samples ($i.e.$, backdoor behaviors) is also widely used [18]. A line of research utilizes backdoor techniques [73, 53, 49, 79, 29] to mark the dataset. Any model trained on the marked dataset will misclassify samples containing specific trigger patterns. Instead, Domain Watermark (DW) [31] introduces EFs that can make the model correctly classify some ‘hard’ samples with high accuracy. DW does not intentionally lead to misclassification and thus is more harmless.
III-C Design Objectives
The objectives of designing a feasible dataset auditing method can be summarized as follows:
  • Effectiveness: The dataset auditing should precisely identify the usage of the dataset from the suspicious model.
  • Fidelity: The dataset auditing method should have a negligible impact on the normal utility of the dataset, $i.e.$, the model trained on the published dataset $\widehat{\mathcal{D}}$ and the raw dataset $\mathcal{D}$ should have minimum difference regarding the primitive task of the dataset.
  • Robustness: The dataset auditing method should not be significantly affected by an adversary’s attacks $i.e.$, still outputs $1$ when the model $\mathcal{M}$ is indeed trained on $\widehat{\mathcal{D}}$.
  • Distinctiveness: The dataset auditing method should not identify other unused datasets as being used for training, $i.e.$, still outputs 00 when $\mathcal{M}$ is not trained on $\mathcal{D}_{u}$.
  • Stealthiness: The perturbations added to the published dataset $\widehat{\mathcal{D}}$ cannot be easily perceived.
  • Efficiency: A good dataset auditing method should introduce as little additional time overhead as possible.
In this paper, we primarily focus on two objectives, robustness and distinctiveness, under adversarial settings.
IV Attack Methodology
As introduced in Section III-C, robustness and distinctiveness are two fundamental and important properties of a practical dataset auditing method. However, there is still a significant gap between the constrained settings of existing works and practice for evaluating these two properties. Existing dataset auditing research [49, 18, 31] primarily considers common yet simple operations in DL and lacks dedicated attack evaluation. To bridge this gap, we propose a unified and comprehensive formulation and methodology of attacks against dataset auditing. We also present systematic attack strategies and implementations of these attacks. All these elements lead to the development of our Dataset Auditing Attack Benchmark, DATABench.
IV-A Attack Formulation
In our threat model, the data owner aims to detect any unauthorized usage of the dataset via dataset auditing. However, an adversary may conduct attacks to obfuscate the results of dataset auditing. In this section, we present two different types of attacks, evasion attack and forgery attack, based on the attack objectives. Intuitively, an evasion attack is to make a ‘trained’ dataset be mistakenly identified as ‘non-trained’. In contrast, a forgery attack is to make a ‘non-trained’ dataset be identified as ‘trained’. Their formal definitions are as follows.
Definition 4 (Evasion Attack against Dataset Auditing).
An evasion attack seeks to conceal the use of a victim dataset. Its goal is to make the dataset auditing method mistakenly infer that the victim dataset was not used during training. Formally, given a dataset $\mathcal{D}$ and any dataset auditing method $\mathcal{A}(\cdot)$, an evasion attack aims to find an approach $\mathcal{E}$ to produce a model $\widetilde{\mathcal{M}}$ trained on $\mathcal{D}$ and have
$\mathcal{A}(\mathcal{D},\widetilde{\mathcal{M}},\mathcal{I})=0,\ \widetilde{\mathcal{M}}=\mathcal{E}(\mathcal{D}),$ (3)
where $\mathcal{I}$ is the auxiliary information required in the process of dataset auditing. The output 00 indicates that $\widetilde{\mathcal{M}}$ is not trained on $\mathcal{D}$. Meanwhile, $\widetilde{\mathcal{M}}$ needs to have a similar utility with the model directly trained on $\mathcal{D}$.
Adversary’s Assumptions in Evasion Attacks. In the evasion attack, the adversary has full control of the model and dataset in the whole training process, including the preprocessing algorithm, the model architecture, and the training method. The adversary also has a small (benign) local dataset related to the model to facilitate evasion attacks.
Definition 5 (Forgery Attack against Dataset Auditing).
A forgery attack aims to falsely attribute the usage of a dataset that was not actually used during training. The adversary attempts to carefully construct a malicious dataset $\mathcal{D}_{a}$ and auxiliary information $\mathcal{I}_{a}$ such that the dataset auditing method will incorrectly identify $\mathcal{D}_{a}$ as being in the training set of an independent model $\mathcal{M}_{u}$, as follows:
$\mathcal{A}(\mathcal{D}_{a},\mathcal{M}_{u},\mathcal{I}_{a})=1,$ (4)
where $\mathcal{M}_{u}$ is not actually trained on $\mathcal{D}_{a}$.
Adversary’s Assumptions in Forgery Attacks. In the forgery attack, the adversary has access to a small clean dataset that is related to the task of the target independent model $\mathcal{M}_{u}$. The adversary also has access to $\mathcal{M}_{u}$. We take two different levels of access into consideration.
  • White-box Access: The adversary has full access to the target model, including the parameters and architecture. This scenario may arise when an adversary attempts to attack an open-source model.
  • Black-box Access: The adversary has API access to the target model, $i.e.$, the inputs and the outputs of the model. This scenario may arise when an adversary tries to attack a model deployed as a service.
IV-B Evasion Attacks
Attack Phases. In evasion attacks, the adversary has full control of the training process and can conduct any operation trying to evade dataset auditing. We split the model training process into three different phases, pre-processing, training, and post-training, as shown in Figure 2.
  1. 1.
    Pre-processing Phase: In this phase, the adversary can apply any pre-process method to the dataset before training to evade dataset auditing.
  2. 2.
    Training Phase: In this phase, the adversary aims to apply different methods to train a high-performance model and, in the meantime, evade dataset auditing.
  3. 3.
    Post-training Phase: In this phase, we assume that the model has been trained and the parameters of the model are fixed. However, the adversary can change the procedure of model inference to evade dataset auditing.
Figure 2: The attack phases in attacking dataset auditing. In the pre-processing phase, the adversary can apply a pre-processing method to the dataset. In the training phase, the adversary utilizes any training technique to get a well-performed and auditing-free model. In the post-training phase, the adversary can alter the inference process of the deployed model to obfuscate dataset auditing.
Attack Strategy. According to Section III-B, existing dataset auditing methods can be classified into internal feature (IF) based methods and external feature (EF) based methods. This taxonomy can facilitate the design of attacks, as follows.
  1. 1.
    Attack Strategy against IF-based Methods: Since internal features are important and indispensable for the utility of the model, the adversary cannot straightforwardly remove them. However, the adversary can undermine the effectiveness of dataset auditing by weakening the association between the model and the training dataset. This strategy is named Decoupling.
  2. 2.
    Attack Strategy against EF-based Methods: To attack EF-based methods, the external features of a dataset do not have a direct correlation with model utility. They primarily lead to the model’s specific behaviors when inputting specific samples. As such, two distinct strategies can be employed to evade EF-based dataset auditing.
    1. (a)
      Removal: This type of attack intends to remove the external features from the dataset and the designated behaviors incurred by the dataset auditing method.
    2. (b)
      Detection: This type of attack aims to detect those specific samples that can trigger the designated unique behaviors of the model and return a random output for them [^4] We consider this type of attack primarily in the post-training phase. This is because it arguably achieves data protection if the adversary detects and removes the marked samples before deploying the model ($i.e.$, the model is not trained on the protected data)..
Attack Implementations. Based on the aforementioned attack strategies, we can design various attacks in different attack phases. The implementations of evasion attacks in DATABench are as follows.
1) Attacks in the Pre-processing Phase.
  • Decoupling: To reduce the association between the trained model and the dataset, the adversary can adopt operations to make the samples in the dataset less similar to the original ones. As such, we design and implement data augmentation and data synthesis as two approaches to achieve such a goal. Data augmentation applies transformations such as cropping and flipping to increase the diversity of training data [78]. Data synthesis is an advanced technique that utilizes the capability of generative models. It first trains a generative model ($e.g.$, a diffusion model [36]) on the dataset and then uses the model to generate a synthetic dataset. The model is trained on the synthetic dataset instead of the original one.
  • Removal: Since the external features usually tend to be ‘noise’ in the dataset, denoising techniques can be adopted to remove them. We consider two classes of denoising methods, traditional denoising filters and DL-based denoising methods. The former reduces noise by smoothing pixel values [24] and the latter exploits DL models ($e.g.$, autoencoder [48]) to remove the noise while preserving the important features for model training.
2) Attacks in the Training Phase.
  • Decoupling: The adversary can reduce the model’s overfitting during training to weaken its association with the training dataset. Regularization is a common approach in practice to achieve such a goal. Additionally, during training, the adversary can apply Differential Privacy (DP) mechanisms [22] ($e.g.$, DP-SGD [1]) to make any single data point indistinguishable.
  • Removal: To eliminate potential abnormal behaviors during training, we consider adversarial training and training-phase backdoor defense as two representative removal attacks. Adversarial training [80] improves the prediction robustness of the model by including adversarial examples (AE) during the training process. AEs are inputs intentionally perturbed to lead to misclassification; Training-phase backdoor defense [26, 40] aims to mitigate hidden external features in the training data by trying not to learn from them. Both ways may remove the designated behaviors and evade dataset auditing.
3) Attacks in the Post-training Phase.
  • Decoupling: In this phase, the adversary can also apply DP mechanisms to add noise during the model inference [19]. As such, we implement noisy output and noisy feature that perturb the outputs and intermediate results of the model, respectively, during inference to evade dataset auditing.
  • Removal: Since the model parameters are fixed, the adversary cannot remove the external features from the model. However, it can still hide the external behaviors by changing inference mechanisms. Randomized smoothing (RS) and transformation-based backdoor defense are two typical methods. Specifically, RS [15] adds noise to the inputs and calculates the aggregated results of the outputs on noisy inputs; Transformation-based backdoor defense [52, 12] ($e.g.$, reprogramming) neutralizes the designated behaviors by applying input transformations.
  • Detection: This type of attack aims to detect the trigger samples used for dataset auditing. Since the trigger samples tend to be different from normal samples, we implement traditional outlier detection based on the features of inputs and advanced backdoor sample detection [37, 30]. If these methods identify the input samples as an outlier or backdoor sample, the adversary can refuse to reply to evade dataset auditing.
IV-C Forgery Attacks
Attack Strategy. The goal of the forgery attack is to craft some samples to falsely claim that an independent model $\mathcal{M}_{u}$ is trained on these samples. Assuming that the adversary has an auxiliary dataset $\mathcal{D}_{a}$, he can optimize a set of perturbations =$\mathcal{R}=\{\bm{r}_{i}\}_{i}^{|\mathcal{D}_{a}|}$ to make the dataset auditing method output $1$. Ideally, this can be formalized as an optimization problem:
$\text{Find}\ \mathcal{R}\ \text{s.t.}\ \mathcal{A}(\mathcal{D}_{a}^{\prime},\mathcal{M}_{u},\mathcal{I}_{a})=1,$ (5)
$\mathcal{D}_{a}^{\prime}$ =$=\left\{(\bm{x}_{i}+\bm{r}_{i},\bm{y}_{i})|(\bm{x}_{i},\bm{y_{i}})\in\mathcal{D}_{a},\bm{r}_{i}\in\mathcal{R}\right\}_{i=1}^{|\mathcal{D}_{a}|},$
where $(\bm{x}_{i},\bm{y}_{i}),\bm{r}_{i}$ are the $i$-th element in $\mathcal{D}_{a}$ and $\mathcal{R}$. Inspired by [56], the adversary can adopt adversarial attack techniques [27, 70] to achieve the forgery attack.
  • IF-based methods primarily rely on the ‘overfitting’ of the internal features to achieve dataset auditing. Therefore, the adversary can increase the output confidence of the model on these forged samples to make them similar to ‘overfitted training samples’.
  • EF-based methods depend on abnormal behaviors (typically, misclassification) of the model. As such, the adversary can optimize the perturbations to lead these forged samples to trigger misclassification.
Attack Implementation. Assuming that $\mathcal{T}(\mathcal{M},\bm{x},\bm{y}_{t})$ is a targeted adversarial attack to achieve the following goal:
$\mathcal{T}(\mathcal{M},\bm{x},\bm{y}_{t})=\arg\min_{\bm{r}}\mathcal{L}(\mathcal{M}(\bm{x}+\bm{r}),\bm{y}_{t}),\ \text{s.t.}\ \|\bm{r}\|\leq\epsilon,$ (6)
and $\bm{y}$ is the ground-truth label of $\bm{x}$. To attack IF-based methods, the adversary can set $\bm{y}_{t}=\bm{y}$ to increase the prediction confidence. In contrast, the adversary can set $\bm{y}_{t}\neq\bm{y}$ to attack EF-based methods. Notably, slight differences exist among the implementations of forged attacks against different dataset auditing methods, but they all follow the same strategy. We include the detailed descriptions in Appendix -D.
Arguably, the adversary can apply any targeted adversarial attacks to achieve such a goal. Specifically, we implement 5 different forgery attacks, as shown in Table I.
TABLE I: A list of attacks against dataset auditing implemented in DATABench. The ‘Data’ column shows whether they need auxiliary data for attacks.
Attack Type Strategy Phase Data
Data Augmentation Evasion Decoupling Pre-processing
Data Synthesis Evasion Decoupling Pre-processing
Gaussian Filter, Median Filter, Wavelet Filter Evasion Removal Pre-processing
Autoencoder-based Denoising Evasion Removal Pre-processing
Regularization, DP-SGD [1] Evasion Decouping Training
Adversarial Training Evasion Removal Training
Adaptively Splitting Poisoned Dataset (ASD) [26] Evasion Removal Training
Noisy Output, Noisy Feature Evasion Decoupling Post-training
Randomized Smoothing (RS) [15] Evasion Removal Post-training
Reprogramming Evasion Removal Post-training
SCALE-UP [30] Evasion Detection Post-training
KNN-based Outlier Detection, SVM-based Outlier Detection Evasion Detection Post-training
FGSM [27], PGD [58], UAP [63] TIFGSM [17], VNIFGSM [82] Forgery / /
TABLE II: Results of evasion attacks in the preprocessing phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack Data Synthesis GaussianFilter MedianFilter WaveletFilter AutoEncoder
IF-based MIA Test Acc. (%, $\uparrow$) 87.58 78.03 87.61 80.38 76.02 85.86
Score ($\uparrow$) 0.62 0.47 0.62 0.55 0.49 0.59
Rapid Test Acc. (%, $\uparrow$) 87.58 78.03 87.61 80.38 76.02 85.86
Score ($\uparrow$) 0.66 0.71 0.66 0.70 0.72 0.65
DI Test Acc. (%, $\uparrow$) 87.58 78.03 87.61 80.38 76.02 85.86
p-value ($\downarrow$) $10^{-25}$ $10^{-3}$ 0.02 0.02 0.36 0.08
Diff ($\uparrow$) 1.33 0.85 0.51 0.51 0.00 0.00
DUA Test Acc. (%, $\uparrow$) 87.17 77.35 87.58 79.31 75.82 84.76
Cost (logits) (%, $\downarrow$) 16.66 100.00 44.11 95.66 22.11 59.60
Cost (label) (%, $\downarrow$) 85.97 100.00 75.80 99.51 22.23 84.14
EF-based DVBW Test Acc. (%, $\uparrow$) 86.58 76.37 86.91 68.48 73.72 83.39
p-value ($\downarrow$) 0.00 0.00 0.00 0.00 1.00 1.00
WSR (%, $\uparrow$) 100.00 99.96 100.00 100.00 19.53 16.40
UBW-P Test Acc. (%, $\uparrow$) 87.28 76.31 87.09 77.25 74.40 84.14
p-value ($\downarrow$) 0.00 $10^{-243}$ 0.00 0.00 1.00 1.00
WSR (%, $\uparrow$) 87.17 39.13 89.17 87.53 29.02 31.16
UBW-C Test Acc. (%, $\uparrow$) 83.05 75.11 82.99 75.79 72.51 81.81
p-value ($\downarrow$) $10^{-60}$ 1.00 $10^{-60}$ 1.00 1.00 $10^{-62}$
WSR (%, $\uparrow$) 75.49 27.61 76.99 33.33 26.86 72.26
ZeroMark Test Acc. (%, $\uparrow$) 87.14 75.37 86.94 70.42 71.66 83.83
p-value ($\downarrow$) $10^{-137}$ $10^{-126}$ $10^{-213}$ 0.96 1.00 1.00
WSR (%, $\uparrow$) 100.00 99.49 100.00 100.00 27.56 19.01
DW Test Acc. (%, $\uparrow$) 86.87 74.88 87.01 77.05 72.45 85.05
WSR (%, $\uparrow$) 67.20 3.70 33.50 85.30 23.00 4.90
V DATABench: Benchmarking Attacks against Dataset Auditing
V-A Overview of DATABench
In this section, we present the overview of our dataset auditing attack benchmark DATABench. Notably, DATABench aims to cover both simple and advanced attack methods. Table I shows the implemented attacks in DATABench. We implement a total of 17 evasion attacks, including 6 attacks in the pre-processing phase, 4 attacks in the training phase, and 7 attacks in the post-training phase. Naive attacks such as pruning and quantization are not included since a wide range of papers have demonstrated they are ineffective [49, 18, 41, 53]. Besides, we implement 5 different forgery attacks. Detailed descriptions of these attacks are in Appendix -B.
In DATABench, we implement 9 representative dataset auditing methods, including 4 IF-based methods and 5 EF-based methods. For IF-based methods, besides Dataset Inference (DI) [60] and Data-use Auditing (DUA) [41], which are designed for dataset auditing, we also adopt two representative MIAs as auditing methods. We select a classic MIA proposed by Shokri et al. [77] (denoted as MIA) and the state-of-the-art (SOTA) MIA named Rapid [35]. For EF-based methods, we evaluate DVBW [53], UBW (including UBW-P and UBW-C) [49], ZeroMark [29], and Domain Watermark (DW) [31]. The detailed settings of these dataset auditing methods can be found in Appendix -A.
TABLE III: Results of evasion attacks in the training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack DP-SGD ($\varepsilon=32$) DP-SGD ($\varepsilon=64$) AdvTraining-FGSM AdvTraining-Hybrid ASD
IF-based MIA Test Acc. (%, $\uparrow$) 87.58 68.54 70.19 83.40 83.44 66.80
Score ($\uparrow$) 0.62 0.48 0.49 0.56 0.53 0.20
Rapid Test Acc. (%, $\uparrow$) 87.58 68.54 70.19 83.40 83.44 66.80
Score ($\uparrow$) 0.66 0.40 0.44 0.69 0.72 0.02
DI Test Acc. (%, $\uparrow$) 87.58 68.54 70.19 83.40 83.44 66.80
p-value ($\downarrow$) $10^{-25}$ 0.06 $10^{-3}$ 0.09 0.18 0.11
Diff ($\uparrow$) 1.33 0.21 0.74 0.19 0.00 0.02
DUA Test Acc. (%, $\uparrow$) 87.17 67.63 70.45 82.98 83.75 79.39
Cost (logits) (%, $\downarrow$) 16.66 100.00 100.00 79.20 99.29 100.00
Cost (label) (%, $\downarrow$) 85.97 100.00 100.00 84.20 98.29 100.00
EF-based DVBW Test Acc. (%, $\uparrow$) 86.58 64.54 66.11 81.44 82.28 77.57
p-value ($\downarrow$) 0.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\uparrow$) 100.00 100.00 100.00 100.00 100.00 78.10
UBW-P Test Acc. (%, $\uparrow$) 87.28 64.95 66.97 81.30 82.22 68.27
p-value ($\downarrow$) 0.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\uparrow$) 87.40 83.90 86.30 87.70 76.80 90.00
UBW-C Test Acc. (%, $\uparrow$) 83.05 66.68 68.67 73.22 70.76 85.38
p-value ($\downarrow$) $10^{-60}$ 1.00 1.00 1.00 1.00 $10^{-41}$
WSR (%, $\uparrow$) 75.49 34.80 35.20 31.10 28.20 44.60
ZeroMark Test Acc. (%, $\uparrow$) 87.14 67.72 69.32 82.17 82.68 78.57
p-value ($\downarrow$) $10^{-137}$ 1.00 1.00 1.00 1.00 1.00
WSR (%, $\uparrow$) 100.00 100.00 100.00 95.70 100.00 99.67
DW Test Acc. (%, $\uparrow$) 86.87 66.12 67.81 85.09 83.91 58.65
WSR (%, $\uparrow$) 67.20 0.00 0.20 2.40 1.20 0.00
TABLE IV: Results of evasion attacks in the post-training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack NoisyOutput NoisyFeature RS Reprogramming SCALE-UP OD-KNN OD-SVM
IF-based MIA Test Acc. (%, $\uparrow$) 87.58 87.26 86.88 81.58 85.21 79.32 84.08 86.57
Score ($\uparrow$) 0.62 0.61 0.61 0.00 0.00 0.50 0.61 0.62
Rapid Test Acc. (%, $\uparrow$) 87.58 87.21 87.30 81.58 85.40 79.32 84.08 86.55
Score ($\uparrow$) 0.66 0.65 0.65 0.19 0.25 0.72 0.70 0.68
DI Test Acc. (%, $\uparrow$) 87.58 87.26 86.88 81.58 85.69 79.32 84.08 86.55
p-value ($\downarrow$) $10^{-25}$ $10^{-5}$ $10^{-9}$ $10^{-7}$ 0.09 $10^{-4}$ 0.36 $10^{-3}$
Diff ($\uparrow$) 1.33 1.15 1.25 1.27 0.24 0.88 0.00 0.70
DUA Test Acc. (%, $\uparrow$) 87.17 87.12 86.93 85.58 85.59 78.46 84.01 86.08
Cost (logits) (%, $\downarrow$) 16.66 31.23 29.97 0.91 51.63 48.37 56.20 37.57
Cost (label) (%, $\downarrow$) 85.97 84.94 70.31 45.77 73.03 63.29 20.43 38.26
EF-based DVBW Test Acc. (%, $\uparrow$) 86.58 86.23 86.13 78.20 83.54 77.91 83.16 85.74
p-value ($\downarrow$) 0.00 0.00 0.00 1.00 1.00 0.00 0.00 1.00
WSR (%, $\uparrow$) 100.00 100.00 100.00 100.00 9.50 81.00 99.63 1.79
UBW-P Test Acc. (%, $\uparrow$) 87.28 86.98 86.98 79.43 75.83 79.40 83.90 86.55
p-value ($\downarrow$) 0.00 0.00 0.00 1.00 1.00 0.00 0.00 0.00
WSR (%, $\uparrow$) 87.39 89.05 89.73 88.29 25.62 87.55 90.00 87.39
UBW-C Test Acc. (%, $\uparrow$) 83.05 82.65 86.32 74.93 84.76 77.49 80.17 82.32
p-value ($\downarrow$) $10^{-60}$ $10^{-97}$ $10^{-97}$ 1.00 1.00 1.00 $10^{-82}$ $10^{-89}$
WSR (%, $\uparrow$) 75.49 77.36 75.74 98.63 76.36 71.76 37.68 77.11
ZeroMark Test Acc. (%, $\uparrow$) 87.14 86.82 86.63 70.45 84.38 76.94 83.97 86.19
p-value ($\downarrow$) $10^{-137}$ 1.00 1.00 1.00 1.00 $10^{-143}$ $10^{-130}$ $10^{-123}$
WSR (%, $\uparrow$) 100.00 100.00 100.00 100.00 10.66 100.00 100.00 100.00
DW Test Acc. (%, $\uparrow$) 86.87 86.79 86.55 82.99 62.97 79.36 82.04 86.08
WSR (%, $\uparrow$) 67.20 61.10 60.60 62.90 11.50 5.60 13.60 62.40
V-B Experimental Settings
Models and Datasets. In the main experiments, we adopt the widely-used CIFAR-10 [46] and ResNet-18 [32]. The reason for using this model and dataset is that it is a simple scenario for dataset auditing, while a challenging one for attacks [18]. Experiments on other models and datasets such as Vision Transformer [61] and ImageNet [16] can be found in Appendix -E. We further split the training data: 70% is the public release dataset (protected), 15% is the defender’s unpublished audit auxiliary dataset, and 15% is the adversary’s unpublished attack auxiliary dataset. Both auxiliary sets are assumed to be trusted and clean (unwatermarked). We train the model from scratch on the victim dataset. The experiments in the fine-tuning scenario are in Appendix -F.
Attack Settings. In our experiments, we evaluate existing dataset auditing methods on all the attacks in DATABench. Given the significant impact of data augmentation and regularization on model utility, we adopt them as default settings. Besides, we consider two variants for adversarial training attacks. ‘AdvTraining-FGSM’ simply uses FGSM [27] to generate the adversarial examples for training, and ‘AdvTraining-Hybrid’ composes three different adversarial attacks [27, 58, 7]. We also test two different privacy budgets ($\varepsilon=32$ or $64$) in DP-SGD to simulate different levels of privacy protection. Detailed descriptions and settings can be found in Appendix -B.
Dataset Auditing Settings. In our experiments, we consider the ideal situation for dataset auditing. If a dataset auditing method needs to train auxiliary models, we assume that the data owner uses the same model architecture as the adversary. It ensures that we evaluate the worst scenario for the adversary. Detailed settings can be found in Appendix -A.
Metrics. We use the accuracy on the testing dataset (Test Acc.) to evaluate the utility of the attacked models. Different metrics are applied to evaluate the effectiveness of different dataset auditing methods, as follows.
  • Score: Score is the average confidence value outputted by the attack models in MIA and Rapid. A higher score indicates that the data is more likely to be a member (trained instance) of the training dataset, and the score $>0.5$ indicates a signal of a ‘trained’ dataset.
  • Diff: Diff is a metric for DI. Diff signifies the average distance between a sample and a neighbouring target class ($i.e.$, decision boundary). A higher Diff indicates that the model is more overfitted to the dataset, suggesting that this dataset is more likely to be a part of the training data.
  • Cost: We utilize Cost (logits) and Cost (label) to reflect the effectiveness of using the logits or only the labels of the model for DUA. Cost (%) is the percentage of samples required for the audit to confidently determine success. A lower cost indicates a more confident judgment, and a $100\%$ cost signifies a signal of ‘non-trained’.
  • WSR: WSR (Watermark Success Rate) measures the percentage of validation data getting the expected prediction, which is defined differently depending on the method: a specific label (DVBW/ZeroMark), an incorrect label (UBW-P/C), or the correct label (DW). We mark WSR $<25\%$ as a signal of a ‘non-trained’ dataset.
  • P-value: The p-value is the output of the hypothesis test used in dataset auditing. The null hypothesis typically assumes that the dataset is independent of the suspicious model, while the alternative hypothesis suggests that they are related. A lower p-value indicates that the null hypothesis is less likely to hold, meaning that the dataset is more likely to be a part of the training dataset. A large number of works depend on validating whether the p-value is lower than $0.05$ to make the decision of auditing.
V-C Results of Evasion Attacks
Tables II-IV show the results of evasion attacks in the pre-processing, training, and post-training phases, respectively. We summarize our key findings as follows.
Takeaway 1: Existing dataset auditing methods are generally vulnerable to evasion attacks.
The experimental results demonstrate that both IF-based and EF-based dataset auditing methods exhibit significant vulnerability when subjected to adversarial evasion attacks across different phases. Across the three tables, numerous orange cells (indicating audit failure) are distributed across various attack phases, covering all evaluated auditing methods. This indicates that the effectiveness of existing dataset auditing techniques is severely challenged when facing adversaries with evasion capabilities. For instance, in the preprocessing phase, the AutoEncoder attack and the Data Synthesis attack completely invalidate 5 and 4 different auditing methods, and WaveletFilter evades 7 of them. In the training phase, several defense mechanisms (like AdvTraining-Hybrid, ASD, DP-SGD) cause multiple IF-based and EF-based methods ($e.g.$, DUA, DI, MIA, Rapid, UBW-C, ZeroMark, and DW under certain attacks) to fail. In the post-training phase, the Reprogramming attack also successfully evades almost all the auditing methods (except DUA).
Takeaway 2: The attack effectiveness targeting a specific class of auditing methods depends on the employed strategy, but cross-class influence is also observed.
The effectiveness of an attack primarily relies on the strategy. For instance, advanced decoupling attacks ($e.g.$, Data Synthesis and DP-SGD) are effective against IF-based methods, and removal attacks ($e.g.$, WaveletFilter and Adversarial Training) work against EF-based methods. In the meantime, cross-class impacts of attacks also exist. WaveletFilter attacks successfully evade DI and MIA. The Reprogramming attack even has a significant negative impact on 8 of 9 methods (including both auditing types). These results demonstrate that attacks primarily designed against one class of methods can also exhibit unintended effectiveness against the other.
Takeaway 3: EF-based methods relying on strong external features show relatively higher robustness but are not immune to evasion attacks.
A key finding is that dataset auditing methods relying on the embedding of strong external features, exemplified by DVBW and UBW-P in our experiments, exhibit relatively higher robustness. Their robustness is particularly exhibited against challenging training-phase attacks like adversarial training and DP-SGD. The two methods are poison-label methods that necessitate changing the labels of part of the samples in the dataset. However, even these methods are not immune to evasion attacks, showing vulnerabilities to specific pre-processing data transformations ($e.g.$, WaveletFilter, AutoEncoder) and post-training inference manipulations ($e.g.$, Reprogramming). Besides, other EF-based methods tested using mild external features ($i.e.$, UBW-C, ZeroMark, DW) displayed considerably less robustness across various attack stages.
TABLE V: Results of the hybrid evasion attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Method Metric No Attack HA1 HA2 HA3
IF-based MIA Test Acc. (%, $\uparrow$) 87.58 78.75 56.93 74.34
Score ($\uparrow$) 0.62 0.00 0.00 0.00
[…]
Score ($\uparrow$) 0.66 0.22 0.16 0.21
DI Test Acc. (%, $\uparrow$) 87.58 79.71 56.74 73.73
p-value ($\downarrow$) $10^{-25}$ 1.00 0.08 0.13
Diff ($\uparrow$) 1.33 0.00 0.00 0.00
DUA Test Acc. (%, $\uparrow$) 87.17 80.39 61.20 73.34
Cost (logits) (%, $\downarrow$) 16.66 99.77 100.00 22.17
Cost (label) (%, $\downarrow$) 85.97 99.86 100.00 23.77
EF-based DVBW Test Acc. (%, $\uparrow$) 86.58 78.10 55.69 70.40
p-value ($\downarrow$) 0.00 1.00 1.00 1.00
WSR (%, $\uparrow$) 100.00 11.47 100.00 19.36
UBW-P Test Acc. (%, $\uparrow$) 87.28 78.51 50.51 71.75
p-value ($\downarrow$) 0.00 1.00 1.00 1.00
WSR (%, $\uparrow$) 87.39 24.82 87.50 28.91
UBW-C Test Acc. (%, $\uparrow$) 83.05 72.81 57.17 70.55
p-value ($\downarrow$) $10^{-122}$ 1.00 1.00 1.00
WSR (%, $\uparrow$) 74.60 35.95 75.49 29.10
ZeroMark Test Acc. (%, $\uparrow$) 87.14 78.78 56.32 75.59
p-value ($\downarrow$) $10^{-137}$ 1.00 1.00 1.00
WSR (%, $\uparrow$) 100.00 13.88 100.00 25.03
DW Test Acc. (%, $\uparrow$) 86.87 77.11 63.61 42.95
WSR (%, $\uparrow$) 67.20 37.20 0.10 6.59
Moreover, IF-based methods generally exhibit moderate robustness, with DUA demonstrating relatively higher robustness. DUA enhances auditing effectiveness by injecting additional noise into the dataset. This mechanism improves robustness but also makes the audit results more susceptible to forgery. We will further discuss the details in Section V-E.
TABLE VI: Results of forgery attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack FGSM PGD UAP TIFGSM VNIFGSM
IF-based MIA Score ($\downarrow$) 0.54 0.66 0.72 0.54 0.71 0.71
Rapid Score ($\downarrow$) 0.22 0.79 0.89 0.30 0.86 0.88
DI p-value ($\uparrow$) 0.01 $10^{-13}$ $10^{-5}$ $10^{-23}$ 0.97 0.01
Diff ($\downarrow$) 0.79 1.35 0.98 1.61 -0.81 0.09
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.13 0.18 0.13 0.24
Cost (label) (%, $\uparrow$) 100.00 0.13 0.13 0.24 0.13 0.24
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 10.01 29.92 98.37 65.60 69.77 96.77
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 11.73 86.55 100.00 71.08 95.69 100.00
ZeroMark p-value ($\uparrow$) 1.00 1.00 1.00 0.99 0.01 1.00
WSR (%, $\downarrow$) 10.01 29.93 98.44 68.28 71.68 96.88
DW WSR (%, $\downarrow$) 11.73 67.00 89.40 66.37 73.72 89.36
V-D Results of Hybrid Evasion Attacks
We also explore combining evasion attack methods from different phases to enhance overall attack effectiveness. Specifically, we hereby evaluate three types of hybrid evasion attacks: (1) HA1: AutoEncoder, Adversarial Training, and Reprogramming. (2) HA2: WaveletFilter, DP-SGD, and Randomized Smoothing. (3) HA3: WaveletFilter and Reprogramming. The results are shown in Table V.
Takeaway 4: Combining attacks from different phases can significantly enhance the attack effectiveness.
Our empirical results corroborate this takeaway, demonstrating the heightened threat posed by hybrid evasion attacks. For instance, HA1 and HA3 successfully circumvent nearly all evaluated auditing methods, substantially degrading the performance metrics and causing audit failures. HA2 also demonstrates considerable efficacy, evading all the auditing methods, but also degrades the utility of the trained models. This finding highlights the feasibility of a sophisticated adversary to combine attacks to achieve better attack effectiveness.
V-E Results of Forgery Attacks
For forgery attacks, we assume that the adversary wants to make their own dataset ($i.e.$, the attack auxiliary dataset) to be falsely identified as the ‘trained dataset’ of other independent models. We use the model trained on the unmarked release dataset as the independent model. The results of white-box forgery attacks are in Table VI, and the experiments in the black-box scenario can be found in Appendix -G.
Figure 3: Relative time overhead of different evasion attacks. The overhead is compared with the baseline time without any specific evasion attack (represented as ‘No Attack’ with a relative overhead normalized to 1.0).
Takeaway 5: Existing dataset auditing methods are overall vulnerable to forgery attacks.
Our experimental results strongly support this finding. When subjected to various forgery attacks, nearly all evaluated auditing methods exhibited significant failures. For instance, IF-based methods, which rely on detecting subtle differences in model behavior (like prediction confidence or loss), were readily deceived. The forgery attacks successfully manipulated model outputs on the forged samples to resemble those of genuinely trained data points, causing these methods to incorrectly flag the unused dataset as ‘trained’ (indicated by drastically reduced costs for DUA, low p-values, and large Diff for DI, and increased scores for MIA/Rapid). Similarly, EF-based methods, including DVBW, UBW, and DW, which look for specific EFs, are also compromised. The forgery attacks effectively generated inputs that triggered the expected model behaviors, leading to low p-values and high Watermark Success Rates (WSR), thus falsely indicating the presence of the EFs and confirming usage.
Takeaway 6: Dataset auditing methods that involve more complex decision-making and auditing processes are more resistant to forgery.
While most methods proved vulnerable, those employing more intricate verification mechanisms demonstrated comparatively better resilience. For example, different from other EF-based methods that directly calculate the accuracy of some specific trigger samples, ZeroMark analyzes the boundary gradients of the closest boundary samples. This sophisticated auditing process makes the auditing results hard to forge. Therefore, although the forgery attack improves the WSR, it only succeeds once against ZeroMark, with the p-values in all other cases remaining close to $1$. Methods relying on simpler signals, like raw confidence scores or basic misclassification, are demonstrably easier to deceive.
V-F Attack Overhead
Beyond evaluating the effectiveness of evasion attacks, understanding their computational overhead is crucial for assessing their practical feasibility for an adversary. An adversary might be constrained by computational resources, potentially limiting their choice of attacks. To quantify this, we measured the execution time required for applying each evasion attack implemented in DATABench during the model training or deployment process, using our primary experimental setup. For each attack, we evaluate the time overhead of the entire pipeline, including data preprocessing, training, and post-training model testing. Figure 3 visualizes the relative time overhead of each attack compared to the baseline time without any specific evasion attack (represented as ‘No Attack’ with a relative overhead normalized to 1.0). The overheads are evaluated using one NVIDIA RTX 3090 GPU.
Our analysis, depicted across the four panels of Figure 3 corresponding to the different attack phases, reveals significant variations in computational cost.
  • In the pre-processing phase, most of the attacks (except data synthesis) exhibit very similar overhead. Data synthesis stands out as exceptionally expensive. It incurs an overhead approximately 65 times greater than baseline. This substantial cost is primarily due to the necessity of training a diffusion model and generating synthesized data using the trained generative model.
  • In the training phase, integrating evasion mechanisms directly into the training loop generally introduces noticeable overhead. Attacks during the training phase incur at least a 1.5× and at most a 13.6× increase in time overhead compared to the baseline.
  • In the post-training phase, all the attacks are consistently lightweight. Most methods demonstrate relative time overheads close to $1.0$. This is expected, as these techniques typically involve minor additions or modifications to the standard inference pipeline, which is computationally much less demanding than model training.
Figure 4: The summary of evaluated dataset auditing methods. This figure overall illustrates the relative performance of the 9 evaluated dataset auditing methods across 5 dimensions.
V-G Summary of Evaluated Dataset Auditing Methods
To provide a holistic comparison of the evaluated dataset auditing methods, we summarize their relative performance across five key dimensions: Robustness, Distinctiveness, Fidelity, Stealthiness, and Efficiency according to the design objectives in Section III-C. Figure 4 represents this comparison using a radar chart, where scores closer to the outer edge (1.0) indicate better performance in that dimension.
Our analysis reveals that no single method excels across all five dimensions. The most critical finding is the pervasive lack of adversarial robustness and distinctiveness, as also analyzed in the previous section. On the other hand, fidelity is typically high, indicating minimal impact on model utility in benign settings. A clear divide exists in stealthiness. IF-based methods inherently excel in stealthiness, while EF-based methods depend on the embedded feature. Efficiency spans a wide spectrum, from computationally cheap approaches to highly demanding ones requiring intensive processing.
In conclusion, the comparative analysis highlights a critical gap: current dataset auditing techniques force difficult trade-offs between essential properties. The widespread vulnerability to evasion attacks and forgery attacks, in particular, highlights a critical future direction.
VI Discussion
VI-A Related Works
Since dataset auditing in DL is an emerging topic, there is only a small number of works surveying this domain. Asswad et al. [2] highlighted the importance of protecting data ownership and reviewed several SOTA techniques in a wide range of data ownership. Chandrasekaran et al. [9] provided a comprehensive study about machine learning governance, including data ownership in machine learning. Recently, Du et al. [18] presented the first systematic SoK regarding dataset auditing, primarily focusing on surveying existing solutions. From the broader perspective of AI copyright protection, the majority of existing work focuses on watermarking techniques for models [68], generated content [69, 88], and traditional multimedia [54]. Lukas et al. [57] present a toolbox to evaluate the robustness of DNN watermarking, which is fundamentally different from dataset auditing.
Compared with existing work, our paper makes the following four distinct contributions to the field. (1) This paper focuses on a specific scenario ($i.e.$, auditing the usage of a dataset) where existing techniques to watermark other objects are not feasible. The dataset is specifically used for training a DL model, and the model is the audit target. (2) This paper aims to provide a novel adversarial angle for evaluating existing dataset auditing methods. (3) Existing works fail to provide a comprehensive benchmark and evaluation regarding the SOTA dataset auditing methods. (4) Existing papers rarely discuss forgery attacks against dataset auditing.
VI-B Dataset Auditing for Generative Models
While dataset auditing has been actively explored for classification models, auditing generative models has only recently started to receive attention [59, 66, 38, 47] and remains largely underdeveloped. Many recent results suggest that even achieving stable results useful for auditing in real-world non-adversarial settings is extremely challenging [20, 21]. This difficulty arises from several inherent characteristics of generative models: (1) They are often trained on massive corpora where any individual dataset constitutes only a minuscule fraction, making both internal and external auditing signals significantly diluted. (2) The outputs of generative models are non-deterministic and vary across sampling seeds, further complicating reliable attribution. (3) Many real-world misuse cases manifest at the output level ($e.g.$, a generated image resembling a copyrighted artwork), where neither internal model access nor query APIs are available ($i.e.$, the box-free setting). In such cases, traditional auditing methods, which rely on model behaviors or access, become ineffective or even inapplicable. While very recent works have achieved partial success in specific settings ($e.g.$, in the fine-tuning scenarios [66, 34, 47]), how to design stable and general-purpose auditing algorithms for generative models is still an open challenge in the community.
Given these challenges, existing auditing methods are not maturely applicable and may yield unstable or misleading results when evaluated under adversarial settings. Therefore, we caution that rushing into adversarial robustness evaluations in such an ill-defined setting without reliable baselines may misrepresent the true difficulty of the problem or misguide the community toward premature conclusions. We leave this as an important direction for future work.
VI-C Potential Future Direction
Future Direction for Dataset Auditing. The demonstrated vulnerability of existing dataset auditing methods underscores an urgent need for fundamentally more robust and distinctive techniques. Future work may focus on developing novel auditing mechanisms resilient to these attacks. (1) For evasion, this could involve exploring IFs that are harder to decouple from the model utility. For EF-based methods, future works may seek EFs that are deeply integrated into the model’s core functionality rather than easily removable or detectable trigger patterns. A certified robust mechanism is also urgently needed. (2) For forgery, our empirical study suggests that complicating the auditing process may lead to better resistance. Consequently, exploring more complex methods, such as multi-bit watermarks or cryptography [25, 75], may be a promising future direction to improve dataset auditing.
Limitations and Future Directions for Evaluating Dataset Auditing. While DATABench provides a comprehensive benchmark, there are still some limitations. First, this paper primarily focuses on dataset auditing regarding image datasets since they are currently the most widely studied and feasible methods. There have been some preliminary works studying auditing datasets of other modalities, such as text [41, 84], point clouds [83], and audio [51]. Providing a comprehensive evaluation of these works may be an important future work. Second, this paper does not cover all possible scenarios of dataset usage in DL. We primarily evaluate the two most important and widely adopted settings: training from scratch (Section V) and fine-tuning (Appendix -F). Third, DATABench does not thoroughly explore finer-grained threat models, such as adversaries with varying levels of capabilities (an adversary with very limited computational resources or data). An exhaustive study of the attacker’s capabilities would be a meaningful direction for future work.
VII Conclusion
This paper presented a comprehensive adversarial evaluation of dataset auditing in DL systems, addressing the critical gap between existing auditing methods and their resilience against dedicated attacks. We introduced a novel internal feature and external feature taxonomy to classify auditing techniques, formally defined evasion and forgery attack objectives, and proposed systematic attack strategies to realize them. These contributions led to the development of DATABench, a benchmark featuring 17 evasion and 5 forgery attacks. DATABench also provides an extensible toolbox to design new dataset auditing methods and attacks. This toolbox can also facilitate the unified evaluation of auditing methods. Our empirical experiments of 9 representative auditing methods against DATABench revealed some critical findings, $e.g.$, none of them demonstrated sufficient robustness or distinctiveness under adversarial settings. These results highlight the significant vulnerability of current dataset auditing methods and underscore the urgent need for developing novel and adversarially resilient dataset auditing methods.
Ethic Considerations
This paper focuses on dataset auditing in Deep Learning (DL). The objective of this paper is to advance the understanding of dataset auditing and evaluate the robustness and distinctiveness of existing methods from an adversarial perspective, which is critical to ensuring the transparency and trustworthiness of dataset auditing methods. Our paper intends to warn the community that existing dataset auditing methods may not be reliable under adversarial settings. All datasets used in our papers are publicly available. We also do not use any personally identifiable information or sensitive data.